# MambaNUT: Nighttime UAV Tracking via Mamba and Adaptive Curriculum Learning

### v1

@@ Abstract
Harnessing low-light enhancement and domain adaptation, nighttime UAV tracking has made substantial strides. However, over-reliance on image enhancement, scarcity of high-quality nighttime data, and neglecting the relationship between daytime and nighttime trackers, which hinders the development of an end-to-end trainable framework. Moreover, current CNN-based trackers have limited receptive fields, leading to suboptimal performance, while ViT-based trackers demand heavy computational resources due to their reliance on the self-attention mechanism. In this paper, we propose a novel pure Mamba-based tracking framework (MambaNUT) that employs a state space model with linear complexity as its backbone, incorporating a single-stream architecture that integrates feature learning and template-search coupling within Vision Mamba. We introduce an adaptive curriculum learning (ACL) approach that dynamically adjusts sampling strategies and loss weights, thereby improving the model’s ability of generalization. Our ACL is composed of two levels of curriculum schedulers: (1) sampling scheduler that transforms the data distribution from imbalanced to balanced, as well as from easier (daytime) to harder (nighttime) samples; (2) loss scheduler that dynamically assigns weights based on data frequency and the IOU. Exhaustive experiments on multiple nighttime UAV tracking benchmarks demonstrate that the proposed MambaNUT achieves state-of-the-art performance while requiring lower computational costs. The code will be available.
Figure 1: Compared to SOTA UAV trackers on NAT2024-1 [13], our MambaNUT sets a new record with 83.3% precision and a speed of 75 FPS, while requiring the lowest computational cost. Note that, the size of the bubbles represents the number of parameters; larger bubbles indicate a higher parameter count.
1 Introduction
00footnotetext: * Corresponding Author: Shuiwang Li.
@@ 1 Introduction
Figure 2: Training data distribution across various datasets, with the sample count varying sharply between daytime and nighttime.
@@
In recent years, many researchers are eager to use low-light image enhancement techniques for nighttime UAV tracking [66, 12, 65, 63]. For example, Fu et al. [12] propose a light enhancer called ”HighlightNet” designed to illuminate specific target areas for UAV trackers. To avoid excessive enhancement in scenarios with complex illumination, LDEnhancer [63] improves nighttime UAV tracking by suppressing light distribution. On the other hand, domain adaptation (DA) is introduced in nighttime UAV tracking, providing an effective solution to the challenges of domain discrepancy. UDAT [67] generate nighttime training samples and adversarially train a model for narrowing the gap between day and night circumstances. TDA-Track [13] proposes a prompt-driven temporal domain adaptation training framework to fully utilize temporal contexts for nighttime UAV tracking. Despite these advancements, current solutions for nighttime tracking continue to face substantial limitations. The development of an end-to-end trainable UAV vision system is hindered by over-reliance on image enhancement, limited availability of high-quality nighttime data, and an often-overlooked relationship between daytime and nighttime trackers. Current CNN-based trackers have limited receptive fields, leading to suboptimal performance, whereas ViT-based trackers demand substantial memory and computational resources due to their reliance on the self-attention mechanism. Additionally, in UAV tracking, the inconsistent feature distribution across consecutive frames hampers long-term object tracking, making long sequence modeling capabilities essential. Recently, the State Space Model has excelled in modeling long-range dependencies with linear complexity, leading to Mamba’s [17] success across visual tasks, particularly in long sequence modeling like video understanding [36, 61] and high-resolution medical image processing [59, 49]. These successful applications inspired us to adapt Mamba for nighttime UAV tracking, leveraging its long-sequence modeling capabilities to learn robust feature representations in low-illuminated scenarios while maintaining lower computational requirements for effective nighttime tracking. Hence, we propose a compact Mamba-based nighttime UAV tracking framework, termed MambaNUT, which adopts a one-stream architecture with a Vision Mamba backbone and a prediction head.
@@
Additionally, class imbalance is an inherent problem in real-world object detection and classification, often causing algorithms to be biased toward the majority classes [30]. In visual tracking, there is a similar imbalance in data distribution between day and night, with more data available during the day. As shown in Fig. 2, compared to current large-scale datasets such as GOT-10K [25], LaSOT [10], and TrackingNet [47], which predominantly consist of daytime images with few or no nighttime images, labeled nighttime data (i.e., SHIFT-Night [52], ExDark [46], and BDD100K-Night [68]) remains relatively scarce. Addressing data imbalance is crucial in this context, as the minority (nighttime) data is the key focus of our work. Training the tracking model with equal weight for samples under varying light conditions may lead to bias toward the majority daytime data and reduced accuracy for the minority nighttime data. Two promising solutions to the imbalanced data learning challenge are resampling [23, 22, 21] and cost-sensitive learning [72, 30, 8]. However, oversampling can lead to overfitting from repeated minority samples, downsampling may discard valuable majority data, and cost-sensitive learning struggles with defining precise costs for samples across different distributions. Curriculum learning (CL) is the learning paradigm inspired by the way humans and animals learn, gradually progressing from easier to more complex samples during training [3, 26]. Inspired by CL, we introduce Adaptive Curriculum Learning (ACL) into our framework to address this issue, based on the following considerations. We aim for the model to first learn appropriate feature representations during the day to enhance its generalization ability, which will improve the learning of more robust feature representations at night. Hence, we propose a dynamic sampling strategy for assigning data weights that emphasizes hard instances, such as nighttime samples, and introduce a novel loss function called Adaptive Data Balance (ADB) Loss, which effectively addresses the data imbalance between daytime and nighttime while enhancing calibration performance. Extensive experiments substantiate the effectiveness of our method and demonstrate that our MambaNUT achieves state-of-the-art performance. As shown in Fig. 1, our method sets a new record with a precision of 83.3, running efficiently at around 75 frames per second (FPS) on the NAT2024-1 [13] and using only 4.1 million parameters, the lowest in comparison. The contributions of our work are summarized as follows:
@@
  • We introduce a simple yet effective Adaptive Curriculum Learning component to address the learning imbalance between daytime and nighttime data, featuring two curriculum schedulers: a dynamic sampling scheduler and a dynamically weighted loss scheduler.
@@
  • Extensive experiments validate that our MambaNUT surpasses state-of-the-art methods on multiple nighttime tracking benchmarks while using fewer parameters and FLOPs.
2 Related work
@@
Nighttime UAV Tracking. Real-world UAV tracking applications encounter considerable challenges in low-illumination nighttime scenarios, as generic trackers are primarily designed for daytime conditions. Recently, low-light enhancement and domain adaptation (DA) have emerged as the two primary methods for improving nighttime UAV tracking performance. In enhancement-based nighttime UAV tracking [66, 12, 65], numerous types of enhancers are proposed to improve image illumination prior to processing by the trackers. Specifically, Li et al. [34] integrate a low-light image enhancer into a CF-based tracker for robust nighttime tracking, while DarkLighter [66] and HighlightNet [12] also develop low-light enhancers to mitigate extreme illumination and emphasize potential objects. However, the limited relationship between low-light image enhancement and UAV tracking leads to suboptimal performance and increased computational costs when enhancers and trackers are integrated in a plug-and-play manner. For DA training-based nighttime UAV tracking [67, 14, 13], trackers utilize domain adaptation to transfer daytime tracking capabilities to nighttime scenarios. For instance, UDAT [67] proposes using a transformer-based bridging layer to align image features from daytime and nighttime domains, thereby transferring somewhat tracking capabilities to the nighttime domain. TDA-Track [13] introduces a novel temporal domain adaptation training framework for nighttime UAV tracking, making it the first to leverage temporal contexts in training nighttime UAV trackers. Unfortunately, DA-based methods incur higher training costs and are limited by the lack of high-quality target domain data for tracking. To build an end-to-end trainable vision system, DCPT [73] introduces a novel architecture that enables robust nighttime UAV tracking by efficiently generating darkness clue prompts without needing a separate enhancer. However, this enhanced tracker burdens resource-limited UAV platforms by adding even more parameters to an already substantial fully transformer-based base tracker, increasing computational resource requirements and hindering efficiency. In our work, we explore the adaptation of Vision Mamba for nighttime UAV tracking for the first time, leveraging its powerful long-sequence modeling capabilities while ensuring computational costs grow linearly for efficient and accurate tracking.
@@
Vision Mamba Models. Unlike traditional structured State Space Models [18], Mamba employs an input-dependent selection mechanism and a hardware-aware parallel algorithm [17], enabling it to model long-range dependency linearly with sequence length. In the field of natural language processing (NLP), it exhibits comparable performance and better efficiency than Transformers in language modeling for long-sequence. Recently, Mamba’s linear complexity in long-range modeling has proven effective and superior across various visual tasks. In classification tasks, Vim [74] and VMamba [45] have shown outstanding performance by building on Mamba’s success, utilizing a bidirectional scanning mechanism and a four-way scanning mechanism, respectively. It also exhibits great potential in high-resolution image tasks, with many notable works proposed in medical image segmentation, including VM-UNet [49] and Swin-UMamba [44]. Subsequently, in the field of video, VideoMamba [36] offers a scalable and efficient solution for comprehensive video understanding, encompassing both short-term and long-term content. MambaTrack [57] explores a Mamba-based learning motion model for multiple object tracking (MOT). In our work, we propose a novel Mamba-based framework for nighttime UAV tracking that incorporates a Adaptive Curriculum Learning (ACL) component to adaptively optimize the sampling strategy and loss weight, enhancing generalization and discrimination in night tracking.
@@
Curriculum learning. The concept of curriculum learning (CL), first proposed in [3], shows that the strategy of learning from easy to hard significantly enhances the generalization of deep models. While these approaches [1, 20, 29] improve convergence speed and local minima quality, pre-determining the order can create inconsistencies between the fixed curriculum and the model being learned. To address this, Kumar et al. [32] proposed the concept of self-paced learning, where the curriculum is constructed dynamically and without supervision to adjust to the learner’s pace. This seminal concept has inspired numerous variations across a range of computer vision applications, including classification [15, 55, 56], action recognition [54], and object [69, 50] / face detection [41, 60]. Despite its efficacy in these domains, the exploration of curriculum learning in the context of visual tracking remains limited. In contrast, our work is the first to explore the integration of Vision Mamba with curriculum learning in a unified framework for nighttime UAV tracking, introducing two levels of curriculum schedulers: one for dynamic sampling and another for dynamically weighted the loss function, where weights are assigned based on data frequency and the IoU.
@@
Figure 3: Overview of the proposed MambaNUT framework. It consists of a Vision Mamba backbone and a tracking head, integrating an adaptive curriculum learning (ACL) approach to dynamically adjust sampling strategies and loss weights during training.
3 Methodology
@@
In this section, we detail the proposed end-to-end tracking framework, termed MambaNUT. First, we begin with the preliminary of state space models (SSM) and the Mamba. Then, we introduce the Adaptive Curriculum Learning (ACL) component for addressing imbalanced data learning problems, which include two-level curriculum schedulers: the sampling scheduler and the loss function scheduler. Last, the overall architecture of the proposed MambaNUT was described in detail, as shown in Fig. 3.
3.1 Preliminary
@@
3.2 Overview
As shown in Fig. 3, our proposed MambaNUT adopts a one-stream framework, which includes a Visison Mamba-based backbone and a tracking head. The framework takes a pair of images as input, namely the template image $Z\in\mathbb{R}^{3\times H_{z}\times W_{z}}$ and the search image $X\in\mathbb{R}^{3\times H_{x}\times W_{x}}$. These images are respectively split and flattened into patch sequences $P\times P$, and the number of patches for $Z$ and $X$ are $P_{z}=H_{z}\times W_{z}/P^{2}$ and $P_{x}=H_{x}\times W_{x}/P^{2}$. The features extracted from the Vision Mamba backbone are input into the prediction head to generate the final tracking results. To enhance the learning of robust feature representations from nighttime samples, we propose a Adaptive Curriculum Learning (ACL) component for the imbalanced data learning problem, which features two-level curriculum schedulers: (1) a sampling scheduler that transforms the data distribution from imbalanced to balanced, as well as from easier (daytime) to harder (nighttime) samples; (2) a data-dependent dynamically weighted loss function that assigns weights based on data frequency and the IOU. The details of the this component will be elaborated in the subsequent subsections.
3.3 Adaptive Curriculum Learning
Sampling is one of the simple and effective methods to deal with imbalanced data learning. Our sampling scheduler is a key element of the Adaptive Curriculum Learning (ACL) component, dynamically adapting the daytime and nighttime data distribution in a batch from imbalanced to balanced throughout the training process. During training, we assign equal sampling weights to all datasets within each epoch; however, for nighttime datasets, their weights are adjusted by dividing by a constant and multiplying by the epoch number, resulting in a smaller initial proportion of nighttime data that gradually increases as training advances. Given a dataset $d$, its assigned sampling weight can be expressed as follows:
@@ 3.3 Adaptive Curriculum Learning
where $e$ refers to current training epoch, $\theta$ represents a constant, set to 150, which is half of the total training epochs. $\mathcal{N}$ are the nighttime datasets. Then, the final sampling ratio for a given dataset among the combination of training sets is: $r_{i}=w_{i}/\sum_{i=1}^{N}w_{i}$, where $N$ denotes the number of training datasets. Usually, the model learns lots of easy (daytime) samples in early stage of the training process. Going further with the training process, the data distribution between daytime and nighttime is gradually getting balanced. During the training phase, the backpropagation algorithm updates the network’s parameters based on the errors computed by the loss function. Training the tracking model with equal weights for samples under varying lighting conditions leads to imbalanced adaptation, caused by the significant distribution disparity between daytime and nighttime, where nighttime images have lower contrast, brightness, and signal-to-noise ratio, causing the tracker to be biased toward daytime conditions. In our work, the minority nighttime samples are the key instances of interest in this learning task.
@@
In view of this, we introduce an Adaptive Data Balanced (ADB) Loss that assigns weights based on the frequency of daytime/nighttime data and IoU, thereby dynamically focusing more on the challenging minority samples, i.e., the nighttime data. For convenience, let the $IOU$ between the predicted boxes and the ground truth be denoted as $U$. Thus, $U_{i}$ is the $IOU$ of the instance $x_{i}$. Inspired by [11], the proposed ADB is formulated as follows:
$\mathcal{L}_{ADB}=-\frac{1}{n}\sum_{i=1}^{n}\omega_{i}^{(1-U_{i})}log(U_{i})-U_{i}(1-U_{i})$ (4)
@@
where $\omega_{i}$ is a hyperparameter determined based on the frequency of data. In the context of classification, $w_{i}$ is typically inversely proportional to the frequency of the classes, allowing it to effectively penalize the majority classes. In our implementation, we define $\omega_{i}$ as the logic ratio of the data frequency of the majority type, defined as: $\omega_{i}=log(N_{max}/N_{j})+0.5]$, where $N_{max}$ denotes the total sample size of the largest training dataset, specifically one of the daytime datasets, and $N_{j}$ represents the total sample size of the dataset to which the $i$-th sample belongs. Adding 0.5 to the log weights to avoid situations where the weight equals zero. If an instance belongs to a dataset with a large number of samples, its weight is relatively small, and vice versa. With this setup, the minority nighttime data contribute more to the network’s gradient calculation, allowing the network to focus less on the majority daytime data and more on the minority during training. $U_{i}(1-U_{i})$ is the regularization term that penalizes overconfident predictions on the target. As the modulating factor, $\omega^{(1-U_{i})}$ directs the network to focus more on samples with lower IoU values.
@@
Table 1: State-of-the-art comparison on the NAT2024-1 [13], NAT2021 [67], and UAVDark135 [35] benchmarks. The top three results are highlighted in red, blue, and green, respectively. Note that the percent symbol (%) is excluded for Prec., Norm.Prec., and Succ. values.
Method Source NAT2024-1[13] NAT2021[67] UAVDark135[35] Avg.FPS FLOP(G) Params.(M)
@@
MambaNUT-Small Ours 83.3 76.9 63.6 70.1 64.6 52.4 70.0 69.3 57.1 72 1.1 4.1
MambaNUT-Tiny 78.6 72.6 60.5 64.4 58.8 47.6 66.2 65.9 54.6 113 0.69 2.6
@@
3.4 Vision Mamba for Tracking

### v2

# MambaNUT: Nighttime UAV Tracking via Mamba-based Adaptive Curriculum Learning

@@ Abstract
Harnessing low-light enhancement and domain adaptation, nighttime UAV tracking has made substantial strides. However, over-reliance on image enhancement, limited high-quality nighttime data, and a lack of integration between daytime and nighttime trackers hinder the development of an end-to-end trainable framework. Additionally, current ViT-based trackers demand heavy computational resources due to their reliance on the self-attention mechanism. In this paper, we propose a novel pure Mamba-based tracking framework (MambaNUT) that employs a state space model with linear complexity as its backbone, incorporating a single-stream architecture that integrates feature learning and template-search coupling within Vision Mamba. We introduce an adaptive curriculum learning (ACL) approach that dynamically adjusts sampling strategies and loss weights, thereby improving the model’s ability of generalization. Our ACL is composed of two levels of curriculum schedulers: (1) sampling scheduler that transforms the data distribution from imbalanced to balanced, as well as from easier (daytime) to harder (nighttime) samples; (2) loss scheduler that dynamically assigns weights based on the size of the training data and IoU of individual instances. Exhaustive experiments on multiple nighttime UAV tracking benchmarks demonstrate that the proposed MambaNUT achieves state-of-the-art performance while requiring lower computational costs. The code will be available at https://github.com/wuyou3474/MambaNUT.
I INTRODUCTION
@@
Figure 1: Compared to SOTA trackers on NAT2024-1 [14], our MambaNUT sets a new record with 83.3% precision and a speed of 75 FPS, while requiring the lowest computational cost. Note that bubble size reflects the number of parameters, with larger bubbles indicating more parameters.
@@
Existing nighttime UAV tracking methods can generally be categorized into two types: trackers based on low-light image enhancement techniques and those utilizing domain adaptation (DA). For the first type of trackers [15, 16, 17], a light enhancer is initially used to brighten the nighttime video, followed by a daytime tracker to locate the object. However, the development of an end-to-end trainable UAV vision system is limited by the reliance on separate enhancement and tracking processes. The latter trackers use domain adaptation to address the domain discrepancy in nighttime UAV tracking [13, 18, 14]. The DA framework consists of a feature generator and a discriminator, where the generator learns to extract domain-invariant features by deceiving the discriminator using both daytime and nighttime training samples. Domain adaptation requires large amounts of data for training, but high-quality target domain samples are scarce for nighttime learning. To address the above issues, DCPT [19] introduces an end-to-end trainable architecture that generates darkness clue prompts to enhance the tracking capabilities of a fixed daytime tracker for nighttime operation, enabling robust nighttime UAV tracking without a separate enhancer. While DCPT demonstrates robust tracking by leveraging critical darkness cues, such ViT-based trackers require significant memory and computational resources because of their reliance on the self-attention mechanism. Recently, the State Space Model has excelled in modeling long-range dependencies with linear complexity, leading to Mamba’s [20] success across visual tasks, particularly in long sequence modeling like video understanding [21] and high-resolution medical image processing [22]. These successful applications inspired us to adapt Mamba for nighttime UAV tracking, leveraging its long-sequence modeling capabilities to learn robust feature representations in low-illuminated scenarios while maintaining lower computational requirements for efficient nighttime tracking. Hence, we propose a compact Mamba-based nighttime UAV tracking framework, termed MambaNUT, which adopts a one-stream architecture with a Vision Mamba backbone and a prediction head.
Figure 2: Training data distribution varies sharply between daytime and nighttime datasets.
@@
Additionally, class imbalance is an inherent problem in real-world object detection and classification, often causing algorithms to be biased toward the majority classes [23]. In visual tracking, there is a similar imbalance in data distribution between day and night, with more data available during the day. As shown in Fig. 2, compared to current large-scale datasets, such as GOT-10K [8], LaSOT [7], and TrackingNet [9], which predominantly consist of daytime images with few or no nighttime images, labeled nighttime data (i.e., SHIFT-Night [24], ExDark [25], and BDD100K-Night [26]) remains relatively scarce. Addressing data imbalance is crucial in this context, as the minority (nighttime) data is the key focus of our work. Two promising solutions to the imbalanced data learning challenge are resampling [27] and cost-sensitive learning [23, 28]. However, oversampling can lead to overfitting from repeated minority samples, downsampling may discard valuable majority data, and cost-sensitive learning struggles with defining precise costs for samples across different distributions. Curriculum learning (CL) is the learning paradigm inspired by the way humans and animals learn, gradually progressing from easier to more complex samples during training [29]. Inspired by CL, we introduce Adaptive Curriculum Learning (ACL) method into our framework to address this issue, based on the following considerations. We aim for the model to first learn appropriate feature representations during the day to enhance its generalization ability, which will improve the learning of more robust feature representations at night. Hence, we propose a dynamic sampling strategy that gradually increases the weight of nighttime data and introduce the Adaptive Data Weighted (ADW) loss, which uses a weighting scheme based on training data size and the IoU of individual instances to effectively emphasize hard cases like nighttime data, enhancing calibration performance. Extensive experiments substantiate the effectiveness of our method and demonstrate that our MambaNUT achieves state-of-the-art performance. As shown in Fig. 1, our method sets a new record with a precision of 83.3%, running efficiently at around 75 frames per second (FPS) on the NAT2024-1 [14] and using only 4.1 million parameters, the lowest in comparison. The contributions of our work are summarized as follows:
@@
  • We introduce a simple yet effective Adaptive Curriculum Learning strategy to address the imbalance learning between daytime and nighttime data, featuring two curriculum schedulers: a dynamic sampling scheduler and a Adaptive Data Weighted (ADW) loss scheduler.
@@
  • Extensive experiments validate that our MambaNUT surpasses state-of-the-art methods on multiple nighttime tracking benchmarks while requiring lower computational costs.
II Related work
@@
Nighttime UAV Tracking. Real-world UAV tracking applications encounter considerable challenges in low-illumination nighttime scenarios, as generic trackers are primarily designed for daytime conditions. Recently, low-light enhancement and domain adaptation (DA) have emerged as the two primary methods for improving nighttime UAV tracking performance. In enhancement-based nighttime UAV tracking [15, 16], numerous types of enhancers are proposed to improve image illumination prior to processing by the trackers. However, the limited relationship between low-light image enhancement and UAV tracking leads to suboptimal performance and increased computational costs when enhancers and trackers are integrated in a plug-and-play manner. For DA training-based nighttime UAV tracking [13, 18, 14], trackers utilize domain adaptation to transfer daytime tracking capabilities to nighttime scenarios. Unfortunately, DA-based methods incur higher training costs and are limited by the lack of high-quality target domain data for tracking. Recently, DCPT [19] introduced a novel architecture that enables robust nighttime UAV tracking by efficiently generating darkness clue prompts to enhance the tracking capabilities of a fixed daytime tracker for nighttime operation, without the need for a separate enhancer, thus developing an end-to-end trainable vision system. However, this enhanced tracker burdens resource-limited UAV platforms by adding even more parameters to an already substantial fully ViT-based base tracker, increasing computational resource requirements and hindering efficiency. In our work, we explore the adaptation of Vision Mamba for nighttime UAV tracking for the first time, leveraging its powerful long-sequence modeling capabilities while ensuring computational costs grow linearly for efficient and accurate tracking.
@@
Vision Mamba Models. Unlike traditional structured State Space Models [30], Mamba employs an input-dependent selection mechanism and a hardware-aware parallel algorithm [20], enabling it to model long-range dependency linearly with sequence length. In the field of natural language processing (NLP), it exhibits comparable performance and better efficiency than Transformers in language modeling for long-sequence. Recently, Mamba’s linear complexity in long-range modeling has proven effective and superior across various visual tasks. In classification tasks, Vim [31] and VMamba [32] have shown outstanding performance by building on Mamba’s success, utilizing a bidirectional scanning mechanism and a four-way scanning mechanism, respectively. It also exhibits great potential in high-resolution image tasks, with many notable works proposed in medical image segmentation, including VM-UNet [22] and Swin-UMamba [33]. Subsequently, in the field of video, VideoMamba [21] offers a scalable and efficient solution for comprehensive video understanding, encompassing both short-term and long-term content. MambaTrack [34] explores a Mamba-based learning motion model for multiple object tracking. In our work, we propose a novel Mamba-based framework for nighttime UAV tracking that incorporates a Adaptive Curriculum Learning (ACL) method to adaptively optimize the sampling strategy and loss weight, enhancing generalization and discrimination in night tracking.
@@
Curriculum learning. The concept of curriculum learning (CL), first proposed in [29], shows that the strategy of learning from easy to hard significantly enhances the generalization of deep models. While these approaches [35, 36, 37] improve convergence speed and local minima quality, pre-determining the order can create inconsistencies between the fixed curriculum and the model being learned. To address this, Kumar et al. [38] proposed the concept of self-paced learning, where the curriculum is constructed dynamically and without supervision to adjust to the learner’s pace. This seminal concept has inspired numerous variations across a range of computer vision applications, including classification  [39], action recognition [40], and object detection [41, 42]. Despite its efficacy in these domains, the exploration of curriculum learning in the context of visual tracking remains limited. In contrast, our work is the first to explore the integration of Vision Mamba with curriculum learning in a unified framework for nighttime UAV tracking, introducing two levels of curriculum schedulers: a dynamic sampling scheduler and a ADW loss scheduler.
@@
Figure 3: Overview of the proposed MambaNUT framework. It includes a Vision Mamba backbone and a tracking head, integrating an adaptive curriculum learning (ACL) approach with two schedulers: (1) a sampling scheduler that balances the data distribution from easier (daytime) to harder (nighttime) samples, and (2) a loss scheduler that assigns weights based on training data size and IoU of individual instances.
III Methodology
@@
In this section, we detail the proposed end-to-end tracking framework, termed MambaNUT. First, we begin with the preliminary of state space models (SSM) and the Mamba. Then, we introduce the Adaptive Curriculum Learning (ACL) strategy for addressing imbalanced data learning problems, which include two schedulers for sampling and loss backward propagation. Last, the overall architecture of the our MambaNUT was described in detail, as shown in Fig. 3.
III-A Preliminary
@@
III-B Overview
As shown in Fig. 3, our proposed MambaNUT adopts a one-stream framework, which includes a Visison Mamba backbone and a tracking head. The framework takes a pair of images as input, namely the template image $Z\in\mathbb{R}^{3\times H_{z}\times W_{z}}$ and the search image $X\in\mathbb{R}^{3\times H_{x}\times W_{x}}$. These images are respectively split and flattened into patch sequences $P\times P$, and the number of patches for $Z$ and $X$ are $P_{z}=H_{z}\times W_{z}/P^{2}$ and $P_{x}=H_{x}\times W_{x}/P^{2}$. The features extracted from the Vision Mamba backbone are input into the prediction head to generate the final tracking results. To enhance the learning of robust feature representations from nighttime samples, we propose a Adaptive Curriculum Learning (ACL) strategy for the imbalanced data learning problem, which features two-level curriculum schedulers: (1) a sampling scheduler that transforms the data distribution from imbalanced to balanced, as well as from easier (daytime) to harder (nighttime) samples; (2) a data-dependent dynamically weighted loss function that assigns weights based on size of training data and the IoU of individual instances. The details of the proposed method will be elaborated in the subsequent subsections.
III-C Adaptive Curriculum Learning (ACL)
Sampling is one of the simple and effective methods to deal with imbalanced data learning. Our sampling scheduler is a key element of the proposed Adaptive Curriculum Learning (ACL) strategy, dynamically adapting the daytime and nighttime data distribution in a batch from imbalanced to balanced throughout the training process. During training, we assign equal sampling weights to all datasets within each epoch; however, for nighttime datasets, their weights are adjusted by dividing by a constant and multiplying by the training epochs, resulting in a smaller initial proportion of nighttime data that gradually increases as training advances. Given a training dataset $d$, its assigned sampling weight can be expressed as follows:
@@ III-C Adaptive Curriculum Learning (ACL)
where $e$ refers to current training epoch, $\theta$ represents a constant, set to 150, which is half of the total training epochs. $\mathcal{N}$ is the set of all nighttime datasets. Then, the final sampling ratio for a given dataset among the combination of training sets is: $r_{i}=w_{i}/\sum_{i=1}^{N}w_{i}$, where $N$ denotes the number of training datasets. Usually, the model learns lots of easy (daytime) samples in early stage of the training process. Going further with the training process, the data distribution between daytime and nighttime is gradually getting balanced. During the training phase, the back propagation algorithm updates the network’s parameters based on the errors computed by the loss function. Training the tracking model with equal weights for samples under varying lighting conditions leads to imbalanced adaptation, caused by the significant distribution disparity between daytime and nighttime, where nighttime images have lower contrast, brightness, and signal-to-noise ratio, causing the tracker to be biased toward daytime conditions. In our work, the minority nighttime samples are the key instances of interest in this learning task.
@@
In view of this, we introduce an Adaptive Data Weighted (ADW) loss that assigns weights based on the size of the training set and IoU, thereby dynamically focusing more on the challenging minority samples, i.e., the nighttime data. For convenience, let the $IoU$ between the predicted boxes and the ground truth be denoted as $U$. Thus, $U_{i}$ is the $IoU$ of the instance $x_{i}$. Inspired by [44], the proposed ADW is formulated as follows:
$\mathcal{L}_{ADW}=-\frac{1}{n}\sum_{i=1}^{n}\omega_{i}^{(1-U_{i})}log(U_{i})-U_{i}(1-U_{i})$ (4)
@@
where $\omega_{i}$ is a hyperparameter determined by the size of training set. In the context of classification, $\omega_{i}$ is typically inversely proportional to the frequency of the classes, allowing it to effectively penalize the majority classes. In our implementation, we define $\omega_{i}$ as the logic ratio of the size of the largest training set, given by $\omega_{i}=\log(N_{max}/N_{j})+0.5$, where $N_{max}$ denotes the total sample size of the largest training dataset (i.e., one of the daytime datasets), and $N_{j}$ represents the total sample size of the dataset to which the $i$-th sample belongs. Adding 0.5 to the log weights to avoid situations where the weight equals zero. If an instance belongs to a dataset with a large number of samples, its weight is relatively small, and vice versa. With this setup, the minority nighttime data contribute more to the network’s gradient calculation, allowing the network to focus less on the majority daytime data and more on the minority during training. $U_{i}(1-U_{i})$ serves as a regularization term that penalizes uncertain predictions, encouraging the model to produce more confident results. As the modulating factor, $\omega^{(1-U_{i})}$ directs the network to focus more on samples with lower IoU values.
@@
TABLE I: State-of-the-art comparison on the NAT2024-1 [14], NAT2021 [13], and UAVDark135 [45] benchmarks. The top two results are highlighted in red and blue, respectively. Note that the percent symbol (%) is excluded for precision (Prec.), normalized precision (Norm.Prec.), and success rate (Succ.) values.
Method Source NAT2024-1[14] NAT2021[13] UAVDark135[45] Avg.FPS FLOPs(GMac) Params.(M)
@@
MambaNUT Ours 83.3 76.9 63.6 70.1 64.6 52.4 70.0 69.3 57.1 72 1.1 4.1
@@
III-D Vision Mamba for Tracking